# Are My Optimized Prompts Compromised? Exploring Vulnerabilities of LLM-based Optimizers

### v1

@@ 4 Related Works
A line of work treats the LLM itself as the engine for searching and refining prompts (Li et al., 2025). Automatic Prompt Engineer (APE) Zhou et al. (2023) proposes generating and selecting candidate instructions with an LLM, while OPRO (Yang et al., 2024) frames “optimization by prompting,” iteratively proposing new prompts based on scored prior attempts. DSPy (Khattab et al., 2023) compiles modular LLM pipelines and automatically optimizes prompts (and demonstrations) to maximize a user-specified metric, yielding self-improving programs rather than fixed prompt templates. TextGrad “backpropagates” Yüksekgönül et al. (2025) natural-language feedback to improve components (including system prompts), Trace formalizes generative optimization over non-differentiable workflows, and SAMMO (Schnabel and Neville, 2024) performs symbolic prompt-program search with multi-objective criteria. For agentic settings, ExpeL (Zhao et al., 2024) accumulates and distills experience into reusable, prompt-level know-how, while PromptAgent (Wang et al., 2024b) casts prompt optimization as strategic planning with MCTS over expert-level prompt states. Trace (Cheng et al., 2024) formalizes trace-based optimization by viewing the entire generation process as a sequence of decision points and refining prompts through structured feedback over these trajectories. Across these methods, the common recipe is LLM-guided candidate generation, evaluation on held-out data or online feedback, and iterative refinement to improve task metrics. Notably, existing works emphasize performance gains and search strategies; to our knowledge they do not systematically analyze safety risks introduced by feedback loops (e.g., poisoned or manipulative feedback), leaving a gap that our work addresses.

### v2

@@ 4 Related Works
A line of work treats the LLM itself as the engine for searching and refining prompts (Li et al., 2025). Automatic Prompt Engineer (APE) Zhou et al. (2023) proposes generating and selecting candidate instructions with an LLM, while OPRO (Yang et al., 2024) frames “optimization by prompting,” iteratively proposing new prompts based on scored prior attempts. DSPy (Khattab et al., 2023) compiles modular LLM pipelines and automatically optimizes prompts (and demonstrations) to maximize a user-specified metric, yielding self-improving programs rather than fixed prompt templates. TextGrad “backpropagates” Yüksekgönül et al. (2025) natural-language feedback to improve components (including system prompts), Trace formalizes generative optimization over non-differentiable workflows, and SAMMO (Schnabel and Neville, 2024) performs symbolic prompt-program search with multi-objective criteria. For agentic settings, ExpeL (Zhao et al., 2024) accumulates and distills experience into reusable, prompt-level insights, while PromptAgent (Wang et al., 2024b) casts prompt optimization as strategic planning with MCTS over expert-level prompt states. Trace (Cheng et al., 2024) formalizes trace-based optimization by viewing the entire generation process as a sequence of decision points and refining prompts through structured feedback over these trajectories. Across these methods, the common recipe is LLM-guided candidate generation, evaluation on held-out data or online feedback, and iterative refinement to improve task metrics. Notably, existing works emphasize performance gains and search strategies; to our knowledge they do not systematically analyze safety risks introduced by feedback loops (e.g., poisoned or manipulative feedback), leaving a gap that our work addresses.
@@ 5 Conclusion
Limitations
Our experiments intentionally fix the backend model to gpt-4.1_2025-04-14 (except for testing the harmscore optimization robustness with gpt-5.1) to isolate the effect of the optimization loop and avoid confounding differences in levels of safety tuning and model capabilities. This choice limits generalizability across model families and deployment settings. We report single-run trends without multiple random seeds or variance estimates due to the computational cost of full prompt-optimization cycles, so stochastic variability is not quantified. We attempt to reduce stochasticity by employing greedy decoding. Finally, our evaluation focuses on HarmBench and WildGuardMix, which primarily measure harmfulness; broader safety dimensions such as truthfulness, bias, and misinformation are not separately covered. Future work should validate these findings across diverse models, multiple seeds, and a wider set of safety benchmarks.
Potential Risks
This work identifies concrete attack vectors against LLM-based prompt optimization pipelines; these insights could be misused to deliberately poison feedback channels or craft inputs that steer optimized system prompts toward unsafe behavior. To reduce misuse risk, we include a lightweight mitigation that practitioners can deploy to harden optimization inputs and feedback boundaries in real systems. The paper may also contain offensive prompts or model outputs as part of safety evaluation; these are included only to support measurement and should be handled with appropriate care. We used icons to warn the reader whenever possible.
Acknowledgement
We used ChatGPT to polish the grammar/writing of this paper.
Appendix A Defense Model for Highlighting
We describe the defense model assumed by the highlighting mitigation.
Defender capabilities.
The defender controls the optimization pipeline structure and can enforce a serialization format for optimizer inputs (e.g., wrapping each query with <query> and </query> tags) before the optimizer processes each batch. The defender does not alter model weights or feedback values.
Defender constraints.
The defender does not inspect, filter, or label individual queries (e.g., due to privacy or scale) and treats query content as opaque. The defense cannot remove attacker content inside a query.
Attacker capabilities.
The attacker can supply or modify a subset of queries and can embed fake feedback tokens inside query text, but cannot modify the pipeline’s enforced formatting or the true feedback channel under this defense setting.
Goal and mechanism.
The goal is to reduce the influence of fake feedback embedded inside queries while preserving the intended optimization objective. The mechanism is purely structural: explicit query delimiters help the optimizer distinguish genuine feedback fields from query content, so poisoned tokens remain confined to the query span.
Appendix B Prompt Evolution of the HarmScore as Feedback Experiment
We provide system prompt snapshots across optimization epochs for the HarmScore as Feedback Experiment.